# Efficient and Accurate Optimal Transport with Mirror Descent and Conjugate Gradients

### v3

@@ A.1 Derivation of dual objective minimization as Bregman projection
$D_{\mathrm{KL}}\big{(}P|P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)\big{)}$ $=\langle P,\log P-\log P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)\rangle+\langle P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)-P,\bm{1}\rangle$
$=\langle P,\log P-{\bm{u}}^{\prime}\bm{1}^{\top}-\bm{1}{{\bm{v}}^{\prime}}^{\top}+\gamma C\rangle+\langle P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)-P,\bm{1}\rangle$
$=\langle P,-\bm{1}+\gamma C+\log P\rangle-\langle{\bm{u}}^{\prime},{\bm{r}}(P)\rangle-\langle{\bm{v}}^{\prime},{\bm{c}}(P)\rangle+\mathrm{const.}$
We shall apply the method of Lagrange multipliers. Given the equality constraints ${\bm{r}}(P)={\bm{r}}$ and ${\bm{c}}(P)={\bm{c}}$, we write the Lagrangian:
${\mathcal{L}}(P,{\bm{u}},{\bm{v}})=\langle P,-\bm{1}+\gamma C+\log P\rangle-\langle{\bm{u}}^{\prime},{\bm{r}}(P)\rangle-\langle{\bm{v}}^{\prime},{\bm{c}}(P)\rangle-\langle{\bm{u}},{\bm{r}}(P)-{\bm{r}}\rangle-\langle{\bm{v}},{\bm{c}}(P)-{\bm{c}}\rangle.$
Taking the partial derivative with respect to $P_{ij}$ yields
$\frac{\partial{\mathcal{L}}}{\partial P_{ij}}=\gamma C_{ij}+\log P_{ij}-u^{\prime}_{i}-u_{i}-v^{\prime}_{j}-v_{j}.$
Setting the partial to zero, we obtain:
$P^{*}_{ij}=\exp(u_{i}+u^{\prime}_{i}+v_{j}+v^{\prime}_{j}-\gamma C_{ij}).$
Plugging the above into the Lagrangian reveals that $-{\mathcal{L}}(P^{*},{\bm{u}},{\bm{v}})=g({\bm{u}},{\bm{v}})$, i.e., minimizing the KL divergence $D_{\mathrm{KL}}\big{(}P|P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)\big{)}$ over the feasible set ${\mathcal{U}}({\bm{r}},{\bm{c}})$ can be cast as minimizing the dual (2.1).
@@
Recall the definition of $\hat{P}^{(t+1)}$ from mirror descent iterates in (4):
@@
$=\frac{1}{\Delta^{(t)}}\langle\nabla h(P^{(t)})-\nabla h(\hat{P}^{t+1}),P^{(t+1)}-P\rangle$ (due to (4))
@@
See 4.1
Proof.
We prove each of the four statements in order.
Proof of the 1st statement. Here, we provide the derivations for (7-8). Consider a mirror descent problem given by (6):
$P^{(t+1)}=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\left\{\langle P,C\rangle+\frac{1}{\Delta_{\gamma}^{(t)}}D_{\mathrm{KL}}(P|P^{(t)})\right\}.$
To derive the dual problem, we write the (scaled) Lagrangian with dual variables ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$ corresponding to equality constraints ${\bm{r}}(P)={\bm{r}}$ and ${\bm{c}}(P)={\bm{c}}$:
${\mathcal{L}}(P,{\bm{u}},{\bm{v}})$ $=\Delta_{\gamma}^{(t)}\langle P,C\rangle+D_{h}(P|P^{(t)})+\langle{\bm{u}},{\bm{r}}-{\bm{r}}(P)\rangle+\langle{\bm{v}},{\bm{c}}-{\bm{c}}(P)\rangle$
$=\Delta_{\gamma}^{(t)}\langle P,C\rangle+\langle P,\log P\rangle-\langle P,\log P^{(t)}\rangle-\langle P-P^{(t)},\bm{1}\rangle+\langle{\bm{u}},{\bm{r}}-{\bm{r}}(P)\rangle+\langle{\bm{v}},{\bm{c}}-{\bm{c}}(P)\rangle.$
Taking the first derivative with respect to the $ij^{\mathrm{th}}$ entry $P_{ij}$ of $P$:
$\frac{\partial{\mathcal{L}}}{\partial P_{ij}}$ $=\Delta_{\gamma}^{(t)}C_{ij}+1+\log P_{ij}-\log P^{(t)}_{ij}-1-u_{i}-v_{j}$
$=\Delta_{\gamma}^{(t)}C_{ij}+\log P_{ij}-\log P^{(t)}_{ij}-u_{i}-v_{j}$
Setting the partial to 00:
$\frac{\partial{\mathcal{L}}}{\partial P_{ij}}=0$ $\iff\log P_{ij}=\log P^{(t)}_{ij}+u_{i}+v_{j}-\Delta_{\gamma}^{(t)}C_{ij}$
$\iff P_{ij}=P^{(t)}_{ij}\exp\{u_{i}+v_{j}-\Delta_{\gamma}^{(t)}C_{ij}\},$
which is equivalent to (7). Under this reparametrization, we have:
$P$ $=P^{(t)}\odot\exp\{{\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\Delta_{\gamma}^{(t)}C\}$
$\log P$ $=\log P^{(t)}+{\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\Delta_{\gamma}^{(t)}C,$
where $\odot$ denotes the Hadamard product.
Now, plugging the above into the Lagrangian and maximizing with respect to the dual variables (and dropping constant terms), we obtain the dual problem (cf. 2.1): {mini}[2] u, vR^n∑_ij P(u, v)_ij - ⟨u, r⟩- ⟨v, c⟩. Let $\hat{{\bm{u}}}^{*}(\gamma^{(t)})$ denote the dual-optimal variables at step $t$ and ${\bm{u}}^{*}(\gamma^{(t)})=\sum_{t=0}^{t}\hat{{\bm{u}}}^{*}(\gamma^{(t)})$ (similarly for ${\bm{v}}$). Unrolling the recursion backwards in time:
$P^{(t)}=P^{(0)}\odot\exp\left\{{\bm{u}}^{*}(\gamma^{(t)})\bm{1}^{\top}+\bm{1}{\bm{v}}^{*}(\gamma^{(t)})^{\top}-\gamma^{(t)}C\right\},$
where we used the definition $\gamma^{(t)}=\sum_{t^{\prime}=0}^{t}\Delta_{\gamma}^{(t^{\prime})}$. Since $P^{(0)}\in\mathbb{R}^{n}_{>0}$ is rank-1 by construction, we can write it in the form $P^{(0)}=\exp\{{\bm{u}}_{0}\bm{1}^{\top}+\bm{1}{\bm{v}}_{0}^{\top}\}$ given some vectors ${\bm{u}}_{0},{\bm{v}}_{0}\in\mathbb{R}^{n}$. That is,
$P^{(t)}=\exp\left\{({\bm{u}}^{*}(\gamma^{(t)})+{\bm{u}}_{0})\bm{1}^{\top}+\bm{1}({\bm{v}}^{*}(\gamma^{(t)})+{\bm{v}}_{0})^{\top}-\gamma^{(t)}C\right\}.$
Then, we observe that problem (A.2) is identical to the dual EOT problem (2.1) up to a simple reparametrization of the dual variables. The result follows since strong duality holds (Slater’s condition is satisfied given the feasible plan ${\bm{r}}{\bm{c}}^{\top}$) and the objective is strictly convex, i.e., the solution $P^{*}(\gamma)$ of the primal is unique.
Proof of the 2nd statement. Observe that at no point in the above proof of the 1st statement did we use the fact that $P^{(t)}\in{\mathcal{U}}({\bm{r}},{\bm{c}})$, nor did we use the optimality of $\hat{{\bm{u}}}^{*}(\gamma^{(t)}),\hat{{\bm{v}}}^{*}(\gamma^{(t)})\in\mathbb{R}^{n}$. As long as each member of the sequence can be written in the form $\widetilde{P}^{(t)}=\widetilde{P}^{(t-1)}\odot\exp\{\hat{{\bm{u}}}(\gamma^{(t)})\bm{1}^{\top}+\bm{1}(\hat{{\bm{v}}}(\gamma^{(t)}))^{\top}-\gamma^{(t)}C\}$ for some dual variables at $\hat{{\bm{u}}},\hat{{\bm{v}}}\in\mathbb{R}^{n}$, the recursion can be unrolled in the same way. That is, the dual problem remains identical in structure (as does its solution), but only the initial values of the dual variables change.
Proof of the 3rd statement. By Lemma A.1, for $P^{(0)}\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ we have
@@
The last inequality holds since $H(P)\geq H({\bm{r}})$ and $H(P)\geq H({\bm{c}})$ for any $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ (Cover, 1999), which together imply $H(P)\geq\max(H({\bm{r}}),H({\bm{c}}))$.
Given the 2nd statement, $P^{(t)}$ remains the same if $P^{(0)}$ was any non-negative rank-1 matrix, which can be written as $\exp\{{\bm{u}}_{0}\bm{1}^{\top}+\bm{1}{\bm{v}}_{0}^{\top}\}$. Hence, the result holds for all such initializations, and not just $P^{(0)}={\bm{r}}{\bm{c}}^{\top}$.
Proof of the 4th statement. First, note that given ${h(P)=\sum_{ij}P_{ij}\log P_{ij}}$, we have ${\nabla h(P)_{ij}=1+\log P_{ij}}$ and ${\nabla h^{-1}(Q)_{ij}=\exp(Q_{ij}-1)}$. Then, given the definition of $\hat{P}^{(t+1)}$ from mirror descent iterates in (4):
@@
$=\langle\nabla f(P^{(t)}),P^{(t)}-P^{(t+1)}\rangle$ (since $f=\langle P,C\rangle$ is linear)
$=\frac{1}{\Delta_{\gamma}^{(t)}}\langle\nabla h(P^{(t)})-\nabla h(\hat{P}^{t+1}),P^{(t)}-P^{(t+1)}\rangle$ (due to (4))
@@
Next, we provide a simple proof for Remark 4.3 See 4.3
@@
Let $B\in{\mathcal{U}}({\bm{r}}^{\prime},{\bm{c}}^{\prime})$ be the transport plan $P({\bm{u}},{\bm{v}})=\exp\{{\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\gamma_{\mathrm{f}}C\}$ after the termination of the main loop (before rounding in L13) of Alg. 1, which takes place after a single outer loop iteration in this setting, since $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$ by construction. Since $B$ is the output of Sinkhorn iteration (Alg. 4), it lies on the simplex, as do its row and column marginals (specifically, we have ${\bm{c}}^{\prime}={\bm{c}}(B)=\tilde{{\bm{c}}}\in\Delta_{n}$ from Alg. 4). Furthermore, $B$ is the unique optimizer of the EOT problem over ${\mathcal{U}}({\bm{r}}^{\prime},{\bm{c}}^{\prime})$ due to Prop. 4.1 and the fact that it has the form $B_{ij}=\exp\{{\bm{u}}_{i}+{\bm{v}}_{j}-\gamma_{\mathrm{f}}C_{ij}\}$:
@@
Sinkhorn iteration used in L7 returns a solution ${\bm{u}},{\bm{v}}$ such that
@@
$\leq\frac{3}{2}\varepsilon^{\prime}+\frac{H_{\min}({\bm{r}}^{\prime},{\bm{c}}^{\prime})}{\gamma_{\mathrm{f}}^{p}}$ (given (25-27))
@@
B An Efficient Line Search Algorithm

### v4

@@ Appendix A Proofs
${\mathcal{L}}(P,\bm{\alpha},\bm{\beta})$ $=\left[D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}+K\right]-\langle\bm{\alpha},{\bm{r}}(P)-{\bm{r}}\rangle-\langle\bm{\beta},{\bm{c}}(P)-{\bm{c}}\rangle$
$=\langle P,-\bm{1}+\gamma C+\log P\rangle-\langle{\bm{u}}+\bm{\alpha},{\bm{r}}(P)-{\bm{r}}\rangle-\langle{\bm{v}}+\bm{\beta},{\bm{c}}(P)-{\bm{c}}\rangle,$ (21)
where we have used $D_{\mathrm{KL}}+K$ as in (20). Recall that the Lagrange dual function is given by (Boyd & Vandenberghe, 2004):
$\inf_{P\in\mathbb{R}^{n\times n}_{>0}}{\mathcal{L}}(P,\bm{\alpha},\bm{\beta}),$
where >$\mathbb{R}^{n\times n}_{>0}$ is the domain of the KL divergence objective in (19). If strong duality holds, we have
$\sup_{\bm{\alpha},\bm{\beta}\in\mathbb{R}^{n}}\inf_{P\in\mathbb{R}^{n\times n}_{>0}}{\mathcal{L}}(P,\bm{\alpha},\bm{\beta})=\min_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\left[D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}+K\right].$ (22)
In this case, we conclude that strong duality holds since Slater’s condition is satisfied, i.e., there exists a strictly feasible $P$; namely, the independence coupling ${\bm{r}}{\bm{c}}^{\top}$. See Ch. 5.2.3 of Boyd & Vandenberghe (2004).
Now, to find the point-wise minimum of ${\mathcal{L}}(P,\bm{\alpha},\bm{\beta})$ with respect to $P$, we require:
$\frac{\partial{\mathcal{L}}}{\partial P_{ij}}=\gamma C_{ij}+\log P_{ij}-u_{i}-\alpha_{i}-v_{j}-\beta_{j}=0.$
Therefore the minimizer $P^{*}$ must have the form
$P^{*}_{ij}=P({\bm{u}}+\bm{\alpha},{\bm{v}}+\bm{\beta};\gamma)=\exp(\alpha_{i}+u_{i}+\beta_{j}+v_{j}-\gamma C_{ij}).$ (23)
for some $\bm{\alpha}$ and $\bm{\beta}$. To eliminate the variable $P$ from $\inf_{P\in\mathbb{R}^{n\times n}_{>0}}{\mathcal{L}}(P,\bm{\alpha},\bm{\beta})$, we plug the form above into the Lagrangian. The first term on the right hand side of (21) is then
$\langle P^{*},-\bm{1}+\gamma C+\log P^{*}\rangle$ $=\langle P^{*},-\bm{1}+\gamma C\rangle+\langle P^{*},({\bm{u}}+\bm{\alpha})\bm{1}^{T}+\bm{1}({\bm{v}}+\bm{\beta})^{T}-\gamma C\rangle,$
$=-\langle P^{*},\bm{1}\rangle+\langle{\bm{u}}+\bm{\alpha},{\bm{r}}(P^{*})\rangle+\langle{\bm{v}}+\bm{\beta},{\bm{c}}(P^{*})\rangle.$
Using this in (21), we find
${\mathcal{L}}(P^{*},\bm{\alpha},\bm{\beta})$ $=-\langle P^{*},\bm{1}\rangle+\langle{\bm{u}}+\bm{\alpha},{\bm{r}}\rangle+\langle{\bm{v}}+\bm{\beta},{\bm{c}}\rangle,$
$=-\sum_{ij}\exp\big{(}(u_{i}+\alpha_{i})+(v_{j}+\beta_{j})-\gamma C_{ij}\big{)}+\langle{\bm{u}}+\bm{\alpha},{\bm{r}}\rangle+\langle{\bm{v}}+\bm{\beta},{\bm{c}}\rangle.$ (24)
where $\bm{\alpha}$ and $\bm{\beta}$ are the only unknowns. Without loss of generality we can change variables in both $P^{*}=P({\bm{u}}+\bm{\alpha},{\bm{v}}+\bm{\beta};\gamma)$ and ${\mathcal{L}}(P^{*},\bm{\alpha},\bm{\beta})$ to $\bm{\alpha}^{\prime}=\bm{\alpha}+{\bm{u}}$ and $\bm{\beta}^{\prime}=\bm{\beta}+{\bm{u}}$, in which case the Lagrangian is equal to
${\mathcal{L}}(P(\bm{\alpha}^{\prime},\bm{\beta}^{\prime};\gamma),\bm{\alpha}^{\prime},\bm{\beta}^{\prime})$ $=-\sum_{ij}\exp(\alpha^{\prime}_{i}+\beta^{\prime}_{j}-\gamma C_{ij})+\langle\bm{\alpha^{\prime}},{\bm{r}}\rangle+\langle\bm{\beta^{\prime}},{\bm{c}}\rangle$
$=-g(\bm{\alpha}^{\prime},\bm{\beta}^{\prime};\gamma),$ (25)
[…]
$\bm{\alpha}^{*},\bm{\beta}^{*}$ $\in\operatorname*{arg\,max}_{\bm{\alpha},\bm{\beta}\in\mathbb{R}^{n}}{\mathcal{L}}(P^{*}(\bm{\alpha},\bm{\beta}),\bm{\alpha},\bm{\beta})$
$=\operatorname*{arg\,min}_{\bm{\alpha},\bm{\beta}\in\mathbb{R}^{n}}g(\bm{\alpha},\bm{\beta};\gamma).$
That is, the solution is identical to the optimal dual solution in (2.1), and therefore Observation A.1 holds.
Lemma 4.1 follows immediately from Observation A.1 using ${P({\bm{u}},{\bm{v}};\gamma)\odot\exp\{-\Delta_{\gamma}C\}=P({\bm{u}},{\bm{v}};\gamma+\Delta_{\gamma})}$. ∎
@@
See 4.2
Proof.
We prove each of the four statements in order.
A.2.1 Proof of the 1st statement.
Here, we first provide a step by step derivation for the following expression for $P^{(t+1)}$ for $t\geq 0$:
$P^{(t+1)}$ $=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}{\bm{u}}^{*}(\gamma^{(t)}),{\bm{v}}^{*}(\gamma^{(t)});\gamma^{(t+1)}\big{)}\Big{)}=P^{*}(\gamma^{(t+1)}),$
where ${\gamma^{(0)}=0}$ by construction, and $P^{*}(\gamma)=P({\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma);\gamma)$ for ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)\in\operatorname*{arg\,min}_{{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}g(\gamma)$, i.e., solutions of the EOT primal (2.1) and dual (2.1) at $\gamma$. Starting with a rank-1 positive matrix $P^{(0)}=\tilde{{\bm{r}}}\tilde{{\bm{c}}}^{\top}=P(\log\tilde{{\bm{r}}},\log\tilde{{\bm{c}}};0)$ given vectors $\tilde{{\bm{r}}},\tilde{{\bm{c}}}$ with positive entries, we obtain using MD updates (6) and Lemma 4.1:
$P^{(1)}$ $=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}\log\tilde{{\bm{r}}},\log\tilde{{\bm{c}}};\Delta_{\gamma}^{(0)}\big{)}\Big{)}$
$=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}\log\tilde{{\bm{r}}},\log\tilde{{\bm{c}}};\gamma^{(1)}\big{)}\Big{)}$ (Since $\gamma^{(0)}=0$ and $\gamma^{(t+1)}=\gamma^{(t)}+\Delta_{\gamma}^{(t)}$ by construction.)
$=P^{*}({\bm{u}}^{*}(\gamma^{(1)}),{\bm{v}}^{*}(\gamma^{(1)});\gamma^{(1)})$ (By Observation A.1.)
$=P^{*}(\gamma^{(1)}).$ (By definition.)
Repeatedly using the same and continuing the iteration:
$\begin{split}P^{(2)}&=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}{\bm{u}}^{*}(\gamma^{(1)}),{\bm{v}}^{*}(\gamma^{(1)});\gamma^{(2)}\big{)}\Big{)}=P^{*}(\gamma^{(2)}),\\ P^{(3)}&=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}{\bm{u}}^{*}(\gamma^{(2)}),{\bm{v}}^{*}(\gamma^{(2)});\gamma^{(3)}\big{)}\Big{)}=P^{*}(\gamma^{(3)}),\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\vdots\\ P^{(t+1)}&=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P\big{(}{\bm{u}}^{*}(\gamma^{(t)}),{\bm{v}}^{*}(\gamma^{(t)});\gamma^{(t+1)}\big{)}\Big{)}=P^{*}(\gamma^{(t+1)}).\end{split}$ (26)
Hence, each iterate $P^{(t+1)}=P^{*}(\gamma^{(t+1)})$ for $t\geq 0$. ∎
A.2.2 Proof of the 2nd statement.
The result follows immediately by applying Observation A.1 to each iterate in (26) to replace ${\bm{u}}^{*}(\gamma),{\bm{v}}^{*}(\gamma)$ terms by arbitrary ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$. ∎
A.2.3 Proof of the 3rd statement.
@@
Recall the definition of $\hat{P}^{(t+1)}$ from mirror descent iterates in (3a):
@@
$=\frac{1}{\Delta^{(t)}}\langle\nabla h(P^{(t)})-\nabla h(\hat{P}^{t+1}),P^{(t+1)}-P\rangle$ (due to (3a))
@@
By Lemma A.2, for $P^{(0)}\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ we have
@@
The last inequality holds since $H(P)\geq H({\bm{r}})$ and $H(P)\geq H({\bm{c}})$ for any $P\in{\mathcal{U}}({\bm{r}},{\bm{c}})$ (Cover, 1999), which together imply $H(P)\geq\max(H({\bm{r}}),H({\bm{c}}))$. ∎
Proof of the 4th statement. First, note that given ${h(P)=\sum_{ij}P_{ij}\log P_{ij}}$, we have ${\nabla h(P)_{ij}=1+\log P_{ij}}$ and ${\nabla h^{-1}(Q)_{ij}=\exp(Q_{ij}-1)}$. Then, given the definition of $\hat{P}^{(t+1)}$ from mirror descent iterates in (3a):
@@
$=\langle\nabla f(P^{(t)}),P^{(t)}-P^{(t+1)}\rangle$ (since $f(P)=\langle P,C\rangle$ is linear, $\nabla_{P}f(P)=C$.)
$=\frac{1}{\Delta_{\gamma}^{(t)}}\langle\nabla h(P^{(t)})-\nabla h(\hat{P}^{t+1}),P^{(t)}-P^{(t+1)}\rangle$ (due to (3a))
@@
Next, we provide a simple proof for Remark 4.4 See 4.4
@@
Let $B\in{\mathcal{U}}({\bm{r}}^{\prime},{\bm{c}}^{\prime})$ be the transport plan $P({\bm{u}},{\bm{v}})=\exp\{{\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\gamma_{\mathrm{f}}C\}$ after the termination of the main loop (before rounding in L13) of Alg. 1, which takes place after a single outer loop iteration in this setting, since $\gamma_{\mathrm{i}}=\gamma_{\mathrm{f}}$ by construction. Since $B$ is the output of Sinkhorn iteration (Alg. 4), it lies on the simplex, as do its row and column marginals (specifically, we have ${\bm{c}}^{\prime}={\bm{c}}(B)=\tilde{{\bm{c}}}\in\Delta_{n}$ from Alg. 4). Furthermore, $B$ is the unique optimizer of the EOT problem over ${\mathcal{U}}({\bm{r}}^{\prime},{\bm{c}}^{\prime})$ due to Prop. 4.2 and the fact that it has the form $B_{ij}=\exp\{{\bm{u}}_{i}+{\bm{v}}_{j}-\gamma_{\mathrm{f}}C_{ij}\}$:
@@
Sinkhorn iteration returns a solution ${\bm{u}},{\bm{v}}$ such that
@@
$\leq\frac{3}{2}\varepsilon^{\prime}+\frac{H_{\min}({\bm{r}}^{\prime},{\bm{c}}^{\prime})}{\gamma_{\mathrm{f}}^{p}}$ (given (32-34))
@@
Appendix B An Efficient Line Search Algorithm
In Section 4.3, we developed the PNCG algorithm, which required a line search procedure. Here, we develop the line search used in our implementation, following a short background on relevant aspects of line search in numerical optimization.